Suppose you have daily returns 0.03990, 0.00429, -0.00045, -0.00544 ... -0.00544, .... for 252 trading days. What is the compound return? How to get it in Excel? Thank you!!!
Answer #1
Daily Return =i
Number of days 252
Compounded Return =R
1+R=(1+i)^252
R=((1+i)^252)-1
i R=((1+i)^252)-1
Daily Return Compounded Return
0.0399 19136.1697
0.00429 1.94105304
-0.00045 -0.107229315
-0.00544 -0.747064907

Luuuuu G12 x fc =((1+F12)^252)-1 2 A B C D E 252 Daily Return =i Number of days Compounded Return =R 1+R=(1+i)^252 R=((1+i)^2